OpenStudy (anonymous):

@jigglypuff314 can you help me out? on 2 questions

OpenStudy (anonymous):

nevermind i send the wrong one hold on

OpenStudy (anonymous):

Part 1: Create a scenario for an arithmetic sequence. For example, Jasmine practices the piano for ______ minutes on Monday. Every day she ___________ her practice time by _________. If she continues this pattern, how many minutes will she practice on the 7th day? Be sure to fill in the blanks with the words that will create an arithmetic sequence. Use your scenario to write the function for the 7th term in your sequence using sequence notation. Part 2: Create a scenario for a geometric sequence. For example, Anthony goes to the gym for ______ minutes on Monday. Every day he _________his gym time by ____________. If he continues this pattern, how many minutes will he spend at the gym on the 5th day? Be sure to fill in the blanks with the words that will create a geometric sequence. Use your scenario to write the formula for the 5th term in your sequence, using sequence notation. Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

i dont understand this

jigglypuff314 (jigglypuff314):

What would you not understand? :) first blank = random number second blank = (increase, decrease, double, half) third blank = small random number then I can help with the rest :P

OpenStudy (anonymous):

the part i part get is Use your scenario to write the function for the 7th term in your sequence using sequence notation.

OpenStudy (anonymous):

no that part i dont get^^

jigglypuff314 (jigglypuff314):

first blank would be y-intercept second blank would be type of slope like positive or negative third blank would be value of slope into slope-intercept form -> y = mx + b

OpenStudy (anonymous):

oh ok thanks!

jigglypuff314 (jigglypuff314):

glad I could help :)

OpenStudy (anonymous):

wait i did 30 min on mondy and 10 days each following day so would it be y=30x+10 or 95=30x+10

jigglypuff314 (jigglypuff314):

y = mx + b where b = y-intercept and m = slope 30 would be the y-intercept and 10 would be the slope :)

jigglypuff314 (jigglypuff314):

so like y = 10x + 30

OpenStudy (anonymous):

and then i dont understand this either; Part 3: Use your scenario from part 2 to write a question that will lead to using the geometric series formula. Use the formula to solve for Sn in your scenario.

OpenStudy (anonymous):

soorry i gotta go, but can you please answer this for me?

jigglypuff314 (jigglypuff314):

yeah, sorry I just realized that there was a difference between an arithmetic sequence and a geometric sequence but I found this http://www.purplemath.com/modules/series3.htm I'll see what I can work out :)

jigglypuff314 (jigglypuff314):

mmm (self-taught on the spot) I think the Sn equation might be \[\large S_n = S(r)^{n+1}\]where S = y-intercept (starting number) r = rate (slope) and n = (number of days) (this is for the geometric sequence) so if you have starts at 30minutes on monday doubles each time from the previous then it would be Sn = (30)(2)^(n+1)

jigglypuff314 (jigglypuff314):

then if it's the 5th day Sn = (30)(2)^(n+1) -> (30)(2)^((5)+1) -> (30)(2)^(6) = 1920 minutes on the 5th day... 0.o